OpenStudy (anonymous):

what is the equivilent fraction to 0.45

OpenStudy (anonymous):

Remember that every decimal number can be re written as a fraction over 10\[.2=\frac{ 2 }{ 10 }=\frac{ 1 }{ 5 }\]\[.02=\frac{ .2 }{ 10 }=\frac{ 2 }{ 100 }=\frac{ 1 }{ 50 }\]and this trend will keep on going, so go ahead and try your problem now and see if you can figure it out now, just remember to simplify whatever fraction you get.

OpenStudy (anonymous):

so 45 over one hundred wood be my fraction to simplyfy ?

OpenStudy (anonymous):

Yes! Excellent work

OpenStudy (anonymous):

thnx:)

OpenStudy (anonymous):

9/20 is final :)

OpenStudy (anonymous):

Completely correct, awesome job